Determine the present value P you must have to invest in the future value A at simple interest rate R after time T.
A=$9500.00 r-%11.0 t-5 years

Answer by stanbon(75887)

Determine the present value P you must have to invest in the future value A at simple interest rate R after time T.
A=$9500.00 r-%11.0 t-5 years
----
A(t) = P(1+r*t)
9500 = P(1+0.11*5)
---
9500 = P(1.55)
---
P = $6129.03
-----------
Cheers,
Stan H.